OpenStudy (anonymous):

can i ask a question in trigo? 1. cos(90degrees-y)=sinx 2. sinsquared A/2=secA-1/2secA 3. cos2A=1-8sinsquared(A/2)cossquared(A/2) 4.sinA=2sin(A/2)cos(A/2) THANKS! :)

OpenStudy (anonymous):

what do you want done?

OpenStudy (anonymous):

are these identities and you need to prove?

OpenStudy (anonymous):

yes. thank you

OpenStudy (anonymous):

for the first one I think you mean 90-x correct?

OpenStudy (anonymous):

no. it's y. my prof gave that question. im so confused

OpenStudy (anonymous):

now, I'm confused... because this I know cos(90-x)=sinx

OpenStudy (anonymous):

do this anyway?

OpenStudy (anonymous):

that's why i'm confused. can you answer the 90-x?

OpenStudy (anonymous):

i'll just ask my prof about it later. but i think you're correct

OpenStudy (anonymous):

here's the first according to what I think the identity is: cos(90 - x) = cos90*cosx + sin90*sinx by the sum/difference rule for cosine =0*cosx + 1*sinx because cos90=0 and sin90 = 1 = sinx QED

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

continue? or are we done?

OpenStudy (anonymous):

i think its done. next question :)

OpenStudy (anonymous):

let's do the easy ones first... #4 next...

OpenStudy (anonymous):

ok. :)

OpenStudy (anonymous):

that's just the double angle formula for sine... let M=A/2 so by the double angle formula for sine, sin(2M) = 2*sinM*cosM now replace M with A/2 so you have sin(2*A/2) = 2*sin(A/2)cos(A/2) sin(A) = 2sin(A/2)cos(A/2)

OpenStudy (anonymous):

the last is the final answer?

OpenStudy (anonymous):

if you mean the end of the proof then, yes.

OpenStudy (anonymous):

thnks my friend. next one? hehehe.

OpenStudy (anonymous):

Let's try #3. btw, did your prof have advice of how to prove identities?

OpenStudy (anonymous):

Like, to prove identities, what do you usually start with, the complicated side or the easier side?

OpenStudy (nottim):

Trix, might want to remove some comments there.

OpenStudy (anonymous):

you can post here, ok? i'll be back tonight. Thank You. GodBless